Lotte Sdn Bhd has incurred RM800,000 researching chemical compounds in the year ended 30 September 2018. It has also spent RM480,000 developing a new product. The product's development was completed on 28 February but management has decided to delay commercial production until July 2019. The product is expected to have a useful life of five years. The development project meets the MFRS 138, Intangible Assets capitalisation criteria.

Required: Explain how should these costs be amortised in the year ended 30 September 2018. [4 marks]
Answer #1

As per MFRS 138

All research costs should not be capitalised instead should be recognised as expense in the year which they were incurred

All development costs should be capitalised and amortised over useful life of Intangible asset

So research costs incurred RM 800,000 for year 30 September 2018 should be recognised as expense in year 30 Sep 2018

So development cost RM 480,000 should be capitalised and should be amortised over life which is 5 years starting for year 30 Sep 2018

Amortisation for one year = RM 480,000/5 = RM 96,000

Amortisation for period 28 Feb 2018 - 30 Sep 2018 = RM 96,000*7/12 = RM 56,000